OpenStudy (anonymous):

An airplane, flying with a tail wind, travels 1300 miles in 2 hours. The return trip, against the wind, takes 2 1/2 hours. Find the cruising speed of the plane and the speed of the wind (assume that both rates are constant).

OpenStudy (anonymous):

do you have an idea of how to solve? ie, the equation you would use?

OpenStudy (anonymous):

I think you have to use elimination but I am honestly lost.

OpenStudy (cwrw238):

let x be cruising speed and y = speed of wind in mph distance = speed * time with tail wind: 1300 = 2(x + y) against wind: 1300 = 2.5(x - y)

OpenStudy (anonymous):

Solve the following for s and w, the plane's speed and wind speed respectively.\[\left\{2(s+w) =1300,\frac{5}{2}(s-w)=1300\right\}\]\[\{s=585 \text{ mph},w=65 \text{ mph}\}\]